# Protecting multimodal large language models against misleading visualizations

### v4

@@ 3.2 Experiment 2 - Assessing MLLM vulnerabilities (Consistency)
Figure 6: Average Likert-scale ratings (1 to 6) on four pairs of misleading and non-misleading visualizations depicting the same data. MLLM results are reported with standard deviations. Human results are reported from Lauer and O’Brien (2020).
Experiment 2 evaluates whether MLLMs interpret consistently two visualizations that represent the same underlying data table, one misleading and the other non-misleading. This is assessed using a Likert-scale question, such as in Figure 2. Unlike in Experiment 1, there is no ground truth answer. Instead, since both visualizations depict the same underlying data, a model is considered misled if it assigns different scores to each of them. We assess the significance of the difference in ratings with a two-sided paired t-test (p-value (p) $\leq$ 0.05).
Figure 6 summarizes the average Likert-scale ratings from all 18 MLLMs for each visualization pair. Human results from prior work (Lauer and O’Brien, 2020) are included for comparison. The average rating difference between the misleading and non-misleading visualizations is similar for humans and MLLMs, though MLLMs tend to assign slightly higher scores overall. For MLLMs, the inconsistency in ratings is statistically significant for the two visualization pairs involving truncated axis: one with bar charts (p = 5e-3) and one with line charts (p = 0.01). For the other two pairs, misrepresentation with bubble charts (p = 0.45) and 3D effects with pie charts (p = 0.55), the differences are not statistically significant. These results further support the conclusion that MLLMs are vulnerable to misleading visualizations, raising questions about their ability to consistently recognize the same underlying data table behind a misleading and a non-misleading version, in particular for truncated axis. However, this vulnerability must be interpreted with nuance, as not all misleaders showed a significant impact on MLLM’s answers consistency.
@@ 3.3 Experiment 3 - Correction methods
Having established that MLLMs are vulnerable to misleading visualizations, we explore six correction methods aimed at improving their robustness. Our focus is on increasing accuracy on the misleading visualizations dataset from Experiment 1, while maintaining performance on the non-misleading dataset. We evaluate the impact of six correction methods using three open-weight, mid-sized MLLMs, Qwen2VL-7B, Ovis1.6-9B, and InternVL2.5-8B, all of which performed below the random baseline on misleading visualizations in Experiment 1. Statistical significance of the changes in accuracy is assessed using the two-sided McNemar test (p $\leq$ 0.05) (McNemar, 1947). Figure 7 shows the change in accuracy relative to Experiment 1 across the six methods: (1) inclusion of a warning message, (2) inclusion of extracted axes, (3) inclusion of the extracted table, (4) inclusion of both axes and table, (5) table-based QA, and (6) redrawing the visualization.
Figure 7: Change ($\Delta$) in accuracy (percentage points) compared to Experiment 1 using different inference-time correction methods. Statistically significant changes ($\leq$0.05) are hashed.
@@
Two correction methods are effective: table-based QA and redrawing the visualization. The most effective approach by far to counter misleaders is table-based QA, which yields significant improvements on misleading visualizations, ranging from 15.4 to 19.6 pp. Its strength lies in fully replacing the visualization with a data table, thereby eliminating misleaders that do not transfer to the tabular format, such as inverted and truncated axis, or misleading annotations. However, this method proves less effective against misleaders that persist even when the data is presented in table form, such as inappropriate item order and concealed uncertainty. Table-based QA comes with a notable cost for non-misleading visualizations. Two models experience significant decreases in accuracy, by up to 8.5 pp. This degradation is primarily due to errors in the intermediate table extraction step, where incorrect values are introduced or relevant information is lost.
@@
Another promising correction method is redrawing the visualization. This method aligns with prior efforts to make human readers more robust to misleading visualizations (Fan et al., 2022). It leads to significant, though more modest, improvements for two MLLMs. The visualization is redrawn using the default settings of the Python library Matplotlib (Hunter, 2007), which inherently avoids certain misleaders like inverting the axes or using inconsistent tick intervals. In Figure 3, the redrawn visualization corrects the misleading trend line by eliminating inconsistent tick intervals, leading to a more accurate representation of the data. This correction method is effective only if the generated code compiles successfully on a Python interpreter; otherwise, the original visualization is used. The lowest redrawing success rates are observed for scatter plots (79%), and stacked bar charts (80%), likely due to their higher visual complexity and number of elements, which require Qwen2.5-7B to produce longer and more detailed Python code. Unlike table-based QA, there are no significant decreases in accuracy on non-misleading visualizations, making it a more attractive option when preserving performance on non-misleading visualization is a priority.
@@
Table extraction is a key intermediate step. Both table-based QA and redrawing the visualization depend heavily on the accuracy of an intermediate step: table extraction. Since the underlying data table is not available, it must be derived from the visualization before either correction method can be applied. This extraction step is often non-trivial, particularly for complex chart types such as maps or scatter plots that contain a high number of visual elements to extract. Additionally, misleaders themselves can degrade MLLMs’ ability to extract data accurately. Perfect accuracy in table extraction is not always necessary, its importance depends on the nature of the question. Some questions can be answered correctly as long as the extracted table preserves the overall trend of the underlying data, while others require exact values. As discussed above, errors in table extraction can significantly decrease accuracy on non-misleading visualizations. In the absence of ground truth tables, we cannot directly evaluate the accuracy of table extraction. Instead, we analyze MLLM outputs on CHARTOM (Rho et al., 2023; Bharti et al., 2024), which pairs each question with two visualizations, one misleading and one non-misleading, generated from the same underlying data. Ideally, the tables extracted from both visualizations should be identical. However, for all MLLMs, the extracted tables have a perfect match in only 4 out of 56 pairs (7.1%), and a partial match for 13 to 14 other pairs (23 to 25%). In Figure 8, despite that the two visualizations represent the same data, the extracted tables show different values and trends. The table extracted from the inverted axis is incorrect. We make the hypothesis that misleaders negatively affect the quality of table extraction too. Given that the two most effective correction methods depend on reliable table extraction, and that extraction errors also negatively impact performance on non-misleading visualizations, improving this intermediate step is a key direction for future work.
@@
Other correction methods are not effective. None of the remaining correction methods yield significant improvements. Including the extracted table in the prompt even results in a significant decrease in accuracy on non-misleading visualizations for one MLLM. While using the extracted table alone proves to be the most effective method, combining it with the original misleading visualization does not produce similar gains. This further suggests that removing or replacing the misleading visualization, rather than merely supplementing it with additional prompt input, is a key factor in the success of a correction method. Although many misleaders rely on axis manipulations, providing the extracted axes alone has little to no effect on performance. Similarly, adding a warning message produces only non-significant changes, with accuracy gains of at most 4.9 pp, mostly limited to the real-world subset. This warning-based approach assumes prior knowledge of the specific misleader in the visualization, making the reported results an upper bound on its effectiveness. In practice, a classifier needs to detect first the presence of misleaders, which remains a challenging task (Lo and Qu, 2025; Alexander et al., 2024). Given the already low results obtained with ground-truth misleader labels, this correction method appears unpromising overall. However, training a highly accurate misleader detection model could enable the selective application of correction methods. This would offer two key advantages: (1) avoiding the application of correction methods to non-misleading visualizations, thereby eliminating any risk of negative impact; and (2) allowing to select the most suitable correction method based on the type of misleader.
@@ 4 Related work
Our correction methods, in particular redrawing the visualization, relate to prior work on visualization linters (Hopkins et al., 2020; Chen et al., 2022; Fan et al., 2022): rule-based methods to detect and correct misleaders. While these tools proved useful to make human readers more robust to misleading visualizations (Fan et al., 2022), they can only operate with specific chart design tools (Hunter, 2007; Hopkins et al., 2020; Chen et al., 2022) or a very limited set of chart and misleader types (Fan et al., 2022). In contrast, our correction methods do not impose such restrictions.
Our findings are further supported by prior work (Bendeck and Stasko, 2025) and parallel studies (Pandey and Ottley, 2025; Chen et al., 2025; Valentim et al., 2025), which also report MLLM vulnerabilities to misleaders (Bendeck and Stasko, 2025; Pandey and Ottley, 2025; Chen et al., 2025) and to other design decisions like the color palette (Valentim et al., 2025). Our study differentiates itself in several key ways. Prior work (Bendeck and Stasko, 2025) only evaluated GPT4 on six misleading visualizations, providing initial hints but lacking depth to fully establish the presence of a vulnerability. Unlike parallel works, we include real-world visualizations and show that vulnerability of MLLMs is somewhat lower in these cases. Furthermore, we provide the first evaluation of six correction methods. Crucially, we examine the trade-off between improving QA performance on misleading visualizations and preserving accuracy on non-misleading ones, an aspect not addressed in other works.
@@ 5 Conclusion
Our findings highlight the vulnerability of MLLMs to misleading visualizations. Experiment 1 demonstrated that MLLMs are unable to accurately interpret the underlying data when presented with misleading visualizations. Experiment 2 showed that MLLMs are particularly susceptible to truncated axes, often failing to provide consistent answers when shown misleading and non-misleading versions of the same data. Identifying this vulnerability fills a critical gap in automated chart understanding research. While MLLMs report strong performance on standard benchmarks such as ChartQA, they remain unprotected against misleading visualizations. This raises serious concerns about their reliability in real-world settings, especially given the potential for such vulnerabilities to be exploited by malicious actors to spread disinformation (Correll and Heer, 2017).
To mitigate this issue, we proposed six inference-time correction methods, two of which, table-based QA and redrawing the visualization, demonstrated significant improvements. However, table-based QA comes at the cost of reduced accuracy on non-misleading visualizations. While most recent work in chart understanding treats visualizations purely as images (Zeng et al., 2025; Zhang et al., 2024; Masry et al., 2025), our findings challenge this direction and sheds a new light on alternative approaches that involve table extraction (Liu et al., 2023a, b; Kim et al., 2025).
6 Limitations
@@
Second, we assume prior knowledge of the chart type (bar, line, …) for generating the prompts for axes extraction and visualization redrawing. This is a reasonable assumption, as the chart type can either be provided by a human user or accurately predicted by a classifier as a preprocessing step.
Third, existing datasets do not provide an equal representation of all misleaders, with truncated and inverted axes being the two most represented categories. This imbalance is prevalent in CALVI (Ge et al., 2023) and present to a lower extent in the real-world cases (Lo et al., 2022).
@@ Appendix A Definition of misleaders
Inverted axis ($n$=26) An axis is oriented in an unconventional direction
@@
Truncated axis ($n$=21) The axis does not start from zero or is truncated
@@
Inappropriate axis range ($n$=15) The axis range is either too broad or too narrow to
@@
Inconsistent tick intervals ($n$=12) Cases with varying intervals between the ticks (Lo et al., 2022).
3D effects ($n$=12) The closer something is, the larger it appears,
@@
Inappropriate item order ($n$=9) The axis labels or legends appear to be in a random
@@
Inappropriate aggregation ($n$=8) Aggregating data in an improper way that leads to
@@
Dual axis ($n$=8) Two independent axes are layered on top of
@@
Misrepresentation ($n$=7) The value labels provided do not match
@@
Cherry picking ($n$=6) Selecting only a subset of data to display, which
@@
Misleading annotations ($n$=5) Annotations that contradict or make it harder
@@
Area encoding ($n$=5) Linearly encoding the values as areas
@@
Concealed uncertainty ($n$=3) Not displaying uncertainty in visualizations may
@@
Missing normalization ($n$=3) Displaying unnormalized data in absolute quantity
@@
Inappropriate use of pie chart ($n$=1) When a pie chart is used for non-part-to-whole data,
@@
Missing data ($n$=1) A visual representation implies data exist but
@@
Overplotting ($n$=1) Displaying too many things on a plot can

### v5

@@ 3.2 Experiment 2 - Assessing MLLM vulnerabilities (Consistency)
Experiment 2 evaluates whether MLLMs interpret consistently two visualizations that represent the same underlying data table, one misleading and the other non-misleading. This is assessed using a Likert-scale question, such as in Figure 2. Unlike in Experiment 1, there is no ground truth answer. Instead, since both visualizations depict the same underlying data, a model is considered misled if it assigns different scores to each of them. We assess the significance of the difference in ratings with a one-sided paired t-test (p-value (p) $\leq$ 0.05).
Figure 5 summarizes the average Likert-scale ratings from all 19 MLLMs for each visualization pair. Average human results from prior work (Lauer and O’Brien, 2020) are included for comparison. The average rating difference between the misleading and non-misleading visualizations is similar for humans and MLLMs, though MLLMs tend to assign slightly higher scores overall. For MLLMs, the inconsistency in ratings is significant for the two visualization pairs involving truncated axis: one with bar charts (p = 3e-3) and one with line charts (p = 4e-3). For the other two pairs, misrepresentation with bubble charts (p = 0.23) and 3D effects with pie charts (p = 0.27), the differences are not significant. These results further support the conclusion that MLLMs are vulnerable to misleading visualizations, raising questions about their ability to consistently recognize the same underlying data table behind a misleading and a non-misleading version, particularly for truncated axis. However, this vulnerability must be interpreted with nuance, as not all misleaders showed a significant impact on MLLM’s consistency.
Figure 5: Average Likert-scale ratings (1 to 6) on four pairs of misleading and non-misleading visualizations depicting the same data. MLLM results are reported with standard deviations. Average human results are reported from Lauer and O’Brien (2020).
@@ 3.3 Experiment 3 - Correction methods
Having established that MLLMs are vulnerable to misleading visualizations, we examine six correction methods designed to enhance their robustness. Our focus is on increasing accuracy on the misleading visualizations dataset from Experiment 1, while maintaining performance on the non-misleading dataset. We evaluate the impact of six correction methods using three open-weight, mid-sized MLLMs, Qwen2VL-7B, Ovis1.6-9B, and InternVL2.5-8B, all of which performed below the random baseline in Experiment 1. Statistical significance of the changes in accuracy is assessed using the two-sided McNemar test (p $\leq$ 0.05) (McNemar, 1947). Figure 6 shows the change in accuracy relative to Experiment 1 across the six methods: (1) inclusion of a warning message, (2) inclusion of extracted axes, (3) inclusion of the extracted table, (4) inclusion of both axes and table, (5) table-based QA, and (6) redrawing the visualization.
@@
Two correction methods are effective: table-based QA and redrawing the visualization. The most effective approach to counter misleaders is table-based QA, which yields significant improvements on misleading visualizations, ranging from 15.4 to 19.6 pp. Originally proposed as a general approach for chart understanding (Liu et al., 2023a, b), table-based QA further shows its effectiveness for countering misleading charts. Its strength lies in replacing the visualization with a data table, thereby eliminating misleaders that do not transfer to the tabular format, such as inverted and truncated axis. However, this method proves less effective against misleaders that persist even when the data is presented in tabular format, such as inappropriate item order. Table-based QA comes with a significant cost for two models on non-misleading visualizations, by up to 8.5 pp. This degradation is due to errors in the intermediate table extraction step, where incorrect values are introduced or relevant information is lost. Appendix E reports table-based QA performance using DePlot (Liu et al., 2023a) and MatCha (Liu et al., 2023b), two methods developed specifically for table extraction.
@@
Another promising correction method is redrawing the visualization. This method aligns with prior efforts to make human readers more robust to misleading visualizations (Fan et al., 2022). It leads to significant, though more modest, improvements for two MLLMs. The visualization is redrawn using the default settings of the Python library Matplotlib (Hunter, 2007), which inherently avoids certain misleaders like inverting the axes or using inconsistent tick intervals. In Figure 3, the redrawn visualization corrects the misleading trend line by eliminating inconsistent tick intervals, leading to a more accurate representation of the data. This correction method is effective only if the generated code compiles successfully on a Python interpreter; otherwise, the original visualization is used. The lowest redrawing success rates are observed for scatter plots (79%) and stacked bar charts (80%), likely due to their higher visual complexity and number of elements, which require Qwen2.5-7B to produce longer Python code. Unlike table-based QA, there are no significant decreases in accuracy on non-misleading visualizations, making it a more attractive option when preserving performance on non-misleading visualizations is a priority.
Appendix F discusses the impact of combining these two effective methods with Chain-of-Thought prompting (Wei et al., 2022).
@@
Table extraction is a key intermediate step. Both table-based QA and redrawing the visualization depend heavily on the accuracy of the intermediate table extraction step. This extraction step is often non-trivial, particularly for maps or scatter plots that contain a high number of visual elements to extract. Additionally, misleaders themselves can degrade MLLMs’ ability to extract data accurately. This is particularly prevalent for inverted axis, 3D effects, and dual axis, as illustrated in Appendix G. Perfect accuracy in table extraction is not always necessary; its importance depends on the nature of the question. Some questions can be answered correctly if the extracted table preserves the overall trend of the underlying data, while others require exact values. We conduct a comprehensive qualitative analysis of table extraction in Appendix H.
Other correction methods are not effective. None of the remaining correction methods yields significant improvements. While using the extracted table alone proves to be the most effective method, combining it with the misleading visualization does not produce similar gains. The MLLMs are biased towards using the image and ignoring the table in most cases.
Although many misleaders rely on axes manipulations, providing the extracted axes has no significant effect on performance. A manual error analysis showed that the axes extraction step is more often accurate than table extraction, as reported in Appendix I. However, unlike the table, the axes alone are insufficient to answer the question. The MLLM still needs to combine the visualization image with the cues from the axes that indicate the presence of misleaders. The results show that this combination is challenging. This further suggests that removing or modifying the misleading visualization, rather than merely supplementing it with additional prompt input, is a key factor in the success of a correction method.
@@
Adding a warning message produces only non-significant changes, with accuracy gains of at most 4.9 pp, mostly limited to the real-world subset. This warning-based approach assumes prior knowledge of the specific misleader in the visualization, making the reported results an upper bound on its effectiveness. In practice, a classifier needs to detect first the presence of misleaders, which remains a challenging task (Lo and Qu, 2025; Alexander et al., 2024). Given the already low results obtained with ground-truth misleader labels, this correction method appears unpromising overall. However, training a highly accurate misleader detection model could enable the selective application of correction methods. This would offer two key advantages: (1) avoiding the application of correction methods to non-misleading visualizations, thereby eliminating any risk of negative impact; and (2) allowing the selection of the most suitable correction method based on the type of misleader.
@@ 4 Related work
The correction methods, particularly misleader warnings and redrawing the visualization, relate to prior work on visualization linters (Hopkins et al., 2020; Chen et al., 2022; Fan et al., 2022): rule-based methods to detect and correct misleaders. While these tools proved useful to make human readers more robust to misleading visualizations (Fan et al., 2022), they can only operate with specific chart design tools (Hunter, 2007; Hopkins et al., 2020; Chen et al., 2022) or a very limited set of chart and misleader types (Fan et al., 2022). In contrast, the correction methods we explored do not impose such restrictions. Table-based QA was originally introduced as a general approach for chart understanding by Liu et al. (2023a, b). Although this approach has been surpassed on standard benchmarks by more recent methods that treat visualizations purely as images Zeng et al. (2025); Masry et al. (2025), our results reveal that table-based QA finds a new purpose as an effective method for counteracting misleading visualizations.
Our findings are further supported by prior work (Bendeck and Stasko, 2025) and parallel studies (Pandey and Ottley, 2025; Chen et al., 2025; Valentim et al., 2025), which also report MLLM vulnerabilities to misleaders (Bendeck and Stasko, 2025; Pandey and Ottley, 2025; Chen et al., 2025) and to other design decisions like the color palette (Valentim et al., 2025). Our study differentiates itself in several key ways. Prior work (Bendeck and Stasko, 2025) only evaluated GPT4 on six misleading visualizations, providing initial hints but lacking depth to fully establish the presence of a vulnerability. Unlike parallel works, we include real-world visualizations and show that the vulnerability of MLLMs is lower in these cases. Furthermore, we provide the first evaluation of six correction methods. Crucially, we examine the trade-off between improving QA performance on misleading visualizations and preserving accuracy on non-misleading ones, an aspect not addressed in other works.
@@ 5 Conclusion
Our findings highlight the vulnerability of MLLMs to misleading visualizations. Identifying this vulnerability fills a critical gap in the research on automated chart understanding. While MLLMs report strong performance on standard benchmarks such as ChartQA, they remain unprotected against misleading visualizations. This raises serious concerns about their reliability in real-world settings, especially given the potential for such vulnerabilities to be exploited by malicious actors to spread disinformation (Correll and Heer, 2017). To mitigate this issue, we evaluated six inference-time correction methods, two of which, table-based QA and redrawing the visualization, demonstrated significant improvements. However, correction methods often come at the cost of reduced accuracy on non-misleading visualizations.
Limitations
@@
Second, we assume prior knowledge of the chart type (e.g., bar, line) for generating prompts for axes extraction and visualization redrawing. This is a reasonable assumption, as the chart type can either be provided by a human user or accurately predicted by a classifier as a preprocessing step.
Third, existing misleading visualization datasets are small compared to standard datasets in the automated chart understanding literature, such as ChartQA. Furthermore, they do not provide an equal representation of all misleaders, with truncated and inverted axes being the two most represented categories. This imbalance is prevalent in CALVI (Ge et al., 2023) and, to a lesser extent, in real-world data. Despite these limitations, CALVI is the reference dataset for assessing human vulnerability to misleading visualizations. It was carefully curated by experts in data visualization and validated on a large sample of human subjects. Therefore, it constitutes a valid resource for evaluating MLLMs. The properties of CALVI are similar to those of VLAT (Lee et al., 2017), an expert-created dataset for non-misleading chart understanding, which acts as a reference despite being much smaller than datasets such as ChartQA. Furthermore, the real-world distribution of misleaders is inherently imbalanced (Lo et al., 2022), with categories such as truncated axis being more prevalent than others. Hence, scaling the real-world dataset would not address the imbalance issues of synthetic datasets, such as CALVI.
Ethics statement
Social impact
Misleading visualizations are a prevalent form of multimodal misinformation. Our results show the importance of protecting not only humans but also MLLMs from misleading visualizations. Our correction methods provides the first promising results in that direction.
Risks
While the proposed correction methods improve the performance of MLLMs on misleading visualizations, they do not guarantee a correct answer. In general, human users should not blindly trust the output of chart understanding systems.
Dataset access
We release the CALVI, VLAT, Likert-scale, and real-world datasets under a CC-BY-SA-4.0 license. The real-world dataset is only intended to be used for research purposes. CHARTOM can be accessed by contacting its authors (Bharti et al., 2024).
AI assistants use
AI assistants were used in this work to assist with writing by correcting grammar mistakes and typos.
@@ Appendix A Definition of misleaders
Table 2 provides the definitions of all the misleaders considered in this work (Lo et al., 2022; Ge et al., 2023), with their number of occurences.
@@
Inverted axis (nn=26) An axis is oriented in an unconventional direction
@@
Truncated axis (nn=21) The axis does not start from zero or is truncated
@@
Inappropriate axis range (nn=15) The axis range is either too broad or too narrow to
@@
Inconsistent tick intervals (nn=12) Cases with varying intervals between the ticks (Lo et al., 2022).
3D effects (nn=12) The closer something is, the larger it appears,
@@
Inappropriate item order (nn=9) The axis labels or legends appear to be in a random
@@
Inappropriate aggregation (nn=8) Aggregating data in an improper way that leads to
@@
Dual axis (nn=8) Two independent axes are layered on top of
@@
Misrepresentation (nn=7) The value labels provided do not match
@@
Cherry picking (nn=6) Selecting only a subset of data to display, which
@@
Misleading annotations (nn=5) Annotations that contradict or make it harder
@@
Area encoding (nn=5) Linearly encoding the values as areas
@@
Concealed uncertainty (nn=3) Not displaying uncertainty in visualizations may
@@
Missing normalization (nn=3) Displaying unnormalized data in absolute quantity
@@
Inappropriate use of pie chart (nn=1) When a pie chart is used for non-part-to-whole data,
@@
Missing data (nn=1) A visual representation implies data exist but
@@
Overplotting (nn=1) Displaying too many things on a plot can